2. An airplane flying parallel to the ground undergoes two consecutive displacements. The first is 75 km 30.0° west of north, and the second is 155 km


1
Expert's answer
2021-01-29T15:45:35-0500

From the condition of the problem the angle between the first and the second direction will be equal to "90\u00b0" . Hence


"d=\sqrt{155^2+75^2}=172.2\ (m)" . Answer


"\tan\alpha=155/75=2.066\to \alpha=64.2°"


The plane moved in the direction "64.2\u00b0-30\u00b0=34.2\u00b0" East of North. Answer